# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v1

@@ 4.1.2 PD vs CTL
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were motor exercises designed for PD assessment, to measure PD’s typical tremor and bradykinesia. We can observe in these tasks that people with PD show significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [13]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but the main set of the metrics is constituted by the higher variations (as standard deviation or amplitudes) of the speed and acceleration on the tablet, a sign of tremor, typical of PD [12]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.
4.1.3 PD vs PDM
Among the 54 task-agnostic metrics, 30 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We present the 10 metrics having the lowest p-values in Table 5 across all tasks. We only find significant metrics for Point and Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can be found in the Point and Spiral tasks and motor exercises designed for PD assessment. We observe few significant metrics for the Point tasks, mostly amplitudes of speed and acceleration, the best being for vertical velocity. PD participants have lower variations in their speed than PDM participants, which are still lower than healthy controls, as we can see in Figure 12. This means we observe less tremor in PDM participants than PD participants. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 13. This also makes the duration of the spiral drawing longer for PDM participants than for others.
Table 5: $p$-values and AUC for the best 10 metrics for PD vs PDM comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
@@ 4.1.3 PD vs PDM
4.2 Task-Specific Results
@@
4.2.1 Point Tasks Results
@@
4.2.2 Spiral Tasks Results
Figure 15 shows the boxplots of the spiral-specific metrics that were significant for at least one of the tasks. All the metrics exhibit a similar behavior across all tasks: PDM and AD participants take more time to draw their spirals than PD and CTL, AD and CTL participants draw more loops, which gives a lower time per loop (Fundamental period of the spiral) for PDM and AD participants. As participants were not instructed to produce a specific number of loops, it is hard to interpret if those findings are related to hyperkinesia in PD participants, finishing faster than the others, or if the decision to draw fewer loops and draw them slower could be linked to a form of cognitive decline.
Table 6: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
@@ 4.2.2 Spiral Tasks Results
Number of Loops 0.370 0.563 $<$0.001 0.615 0.209 0.542
Fundamental Period $<$0.001 0.694 0.109 0.545 0.004 0.615
Dominant Duration $<$0.001 0.695 0.496 0.543 0.155 0.623
Fundamental Period $<$0.001 0.694 0.474 0.545 0.185 0.615
Non-Dominant N loops 0.714 0.573 0.018 0.604 0.166 0.538
Duration 0.123 0.695 0.043 0.543 0.028 0.623
Number of Loops 0.674 0.563 0.031 0.615 0.241 0.542
Fundamental Period 0.100 0.694 0.051 0.545 0.035 0.615
Pataka N loops 0.150 0.573 0.012 0.604 0.429 0.538
Duration $<$0.001 0.695 0.700 0.543 0.105 0.623
Number of Loops 0.406 0.563 0.007 0.615 0.620 0.542
Fundamental Period $<$0.001 0.694 0.658 0.545 0.103 0.615
@@
4.2.3 Numbers Task Results
Table 7: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
@@ 4.2.3 Numbers Task Results
Table 7 shows the Numbers-specific metrics’ results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.
4.2.4 Writing Tasks Results
@@
In this section, we present the results for the writing tasks’s specific metrics. As shown in Figure 16, AD participants take a much longer Average Time Between Writing than all other participants, across all tasks, revealing a longer reflexion time between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
Table 8: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
@@
4.3 Classification Results
This subsection discusses the classification results obtained from the different handwriting tasks analyzed in this study. The results are discussed by examining each group comparison individually. Specifically, the following six group comparisons are analyzed: AD vs. CTL, PD vs. PDM, AD vs. CTL, AD vs. PDM, PD vs. CTL, and CTL vs. PDM. The models’ performance is evaluated in terms of accuracy (ACC), F-1 score (F1), and AUC. Table 9 presents results obtained using the Bagging classifier. The Appendix includes comprehensive experimental results from the RF and MLP classifiers.
Table 9: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
@@ 4.3 Classification Results
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [36]. This task might exacerbate the difficulties encountered by PD subjects when performing the task, as they are required to use their non-dominant hand.
@@
For the comparison between AD and CTL, the Draw Clock task achieves the highest accuracy (87) and F1 score (87). This task assesses the ability to interpret and depict clock hands, a skill often compromised in individuals with early-stage dementia. The high discriminability of this task could be attributed to cognitive impairment present in the AD group, as indicated by their low MoCA score. Additionally, the Copy Text and Point Sustained tasks yield the highest AUC, indicating their utility in distinguishing between AD and CTL groups.
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. This task requires that subjects maintain a pen vertically above a point while sustaining a vowel, demanding sustained attention and concentration. Given the attentional deficits commonly observed in AD, this task’s effectiveness in discriminating between AD and CTL and PDM groups is plausible.
@@ 5 Conclusion
The use of the proposed set of metrics for binary classification yielded interesting results for all the groups of tasks. Both Drawing tasks and Point tasks are very discriminative for AD participants (up to 91% Accuracy against the PD group, up to 87% Accuracy against the CTL group). Meanwhile, the Spiral tasks are more effective in discriminating PDM participants, in particular the Spiral PaTaKa task that shows 73% accuracy against PD, 73% against AD and 77% against CTL. Across all tasks, AD discrimination against CTL is almost always more accurate than PD against CTL, which confirm our hypothesis that handwriting in general contains valuable informations for AD assessement and characterization. In general, the variations of performances within a task group are non-negligeable, which is supprising given that the differences between similar tasks explored in the section 3.5.2 did not show any significance, except for Copy Text vs Copy Read Text, which mean those differences should be studied further, using more generalist systems to tackle groups of tasks at a time.
@@
The limitations observed in our approach suggest that generic solutions may not be fully suited to this specific application. Future research will focus on developing handcrafted machine learning models, tailored and trained to better capture the complexities of handwriting associated with neurodegenerative disease characterization. Additionally, given the multimodal and synchronized nature of the presented dataset, future efforts will explore the potential benefits of integrating all available modalities to enhance performance on the task.

### v2

@@ IV-A2 PD vs CTL
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were better suited to assess motoric features of PD, such as tremor and bradykinesia. Accordingly, we observed subjects with PD showed significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [44]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but metrics of variation (as standard deviation or amplitudes) of the speed and acceleration on the tablet were notable. These findings may represent, tremor, dystonia, dyskinesia, myoclonus other other complicating motor symptoms that develop in some persons with PD [11]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.
IV-A3 PD vs PDM
Among the 54 task-agnostic metrics, 30 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We present the 10 metrics having the lowest p-values in Table V across all tasks. We only find significant metrics for Point and Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can be found in the Point and Spiral tasks and motor exercises designed for PD assessment. We observe few significant metrics for the Point tasks, mostly amplitudes of speed and acceleration, the best being for vertical velocity. PD participants have lower variations in their speed than PDM participants, which are still lower than healthy controls, as we can see in Figure 12. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 13.
TABLE V: $p$-values and AUC for the best 10 metrics for PD vs PDM comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
@@ IV-A3 PD vs PDM
IV-B Task-Specific Results
@@
IV-B1 Point Tasks Results
@@
IV-B2 Spiral Tasks Results
Figure 15 shows the boxplots of the spiral-specific metrics that were significant for at least one of the tasks. All the metrics exhibit a similar behavior across all tasks: PDM and AD participants take more time to draw their spirals than PD and CTL, AD and CTL participants draw more loops, which gives a lower time per loop (Fundamental period of the spiral) for PDM and AD participants. Subjects were not instructed to provide a minimum number of loops, therefore interpretation of task duration is difficult. The reduced number of loops in the PD cohort may have been an energetically rationale choice as the task is particularly difficult with persons with impaired dexterity. As PDM represents a mixed disease cohort, few conclusions relative to PD can be drawn. It is doubtful differences are due to cognitive impairment as overall MOCA performance was similar between these groups.
TABLE VI: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
@@ IV-B2 Spiral Tasks Results
N Loops 0.370 0.563 $<$0.001 0.615 0.209 0.542
Fund. Period $<$0.001 0.694 0.109 0.545 0.004 0.615
Spiral Duration $<$0.001 0.695 0.496 0.543 0.155 0.623
Dom. Fund. Period $<$0.001 0.694 0.474 0.545 0.185 0.615
Spiral N loops 0.714 0.573 0.018 0.604 0.166 0.538
Non Duration 0.123 0.695 0.043 0.543 0.028 0.623
Dom. N Loops 0.674 0.563 0.031 0.615 0.241 0.542
Fund. Period 0.100 0.694 0.051 0.545 0.035 0.615
N loops 0.150 0.573 0.012 0.604 0.429 0.538
Spiral Duration $<$0.001 0.695 0.700 0.543 0.105 0.623
Pataka N Loops 0.406 0.563 0.007 0.615 0.620 0.542
Fund. Period $<$0.001 0.694 0.658 0.545 0.103 0.615
@@
IV-B3 Numbers Task Results
TABLE VII: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
@@ IV-B3 Numbers Task Results
Table VII shows the Numbers-specific metrics’ results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.
IV-B4 Writing Tasks Results
@@
In this section, we present the results for the writing tasks’s specific metrics. As shown in Figure 16, AD participants take a much longer Average Time Between Writing than all other participants across all tasks, revealing a longer hesitation and intermitent pausing times, between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
TABLE VIII: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
@@
IV-C Classification Results
This subsection discusses the classification results obtained from the different handwriting tasks analyzed in this study. The results are discussed by examining each group comparison individually. The models’ performance is evaluated regarding accuracy (ACC), F-1 score (F1), and AUC. Table IX presents results obtained using the Bagging classifier. The Appendix includes comprehensive experimental results from the RF and MLP classifiers.
TABLE IX: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
@@ IV-C Classification Results
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, the latter of which is particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [45]. This is intriguing as mirror movement or mirroring dystonia have been described in Task Specific Dystonia and other movement disorders as a way to unmask and characterize latent dystonia features [46]; in our PD cohort, this task may be abnormal in PD subjects who are prone to the development of dystonic features - a finding generally though to be absent in AD.
@@
For the comparison between AD and CTL, the Draw Clock task achieves the highest accuracy (87) and F1 score (87). This task assesses the ability to draw a clock with two hands correctly placed. This complex task includes a wide set of skills, including adequate function of spatial awareness, executive planning, error detection, motor coordination and execution as well as language processing and numerical processing. This task is often used to screen for intra-cortical and subcortical brain network deficiencies, observed notably in individuals with early-stage dementia. The high discriminability of this task could be attributed to cognitive impairment present in the AD group, as indicated by their lower MoCA score compared to other groups. Additionally, the Copy Text and Point Sustained tasks yield the highest AUC, indicating their utility in distinguishing between AD and CTL groups.
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. Given the attentional deficits commonly observed in AD, this task’s effectiveness in discriminating between AD and CTL and PDM groups is plausible.
@@ V Discussion and Conclusion
The use of the proposed set of metrics for binary classification yielded interesting results for all the groups of tasks. Both Drawing tasks and Point tasks are very discriminative for AD participants (up to 91% Accuracy against the PD group, up to 87% Accuracy against the CTL group). Meanwhile, the Spiral tasks are more effective in discriminating PDM participants, in particular the Spiral PaTaKa task that shows 73% accuracy against PD, 73% against AD and 77% against CTL. Across all tasks, AD discrimination against CTL is almost always more accurate than PD against CTL, which confirm our hypothesis that handwriting in general contains valuable informations for AD assessement and characterization. In general, the variations of performances within a task group are non-negligeable, which is supprising given that the differences between similar tasks explored in the section III-E2 did not show any significance, except for Copy Text vs Copy Read Text, which mean those differences should be studied further, using more generalist systems to tackle groups of tasks at a time.
@@
However, several limitations should be addressed. The sample size restricts generalization, especially considering the age difference between the PDM group and the others. The lack of significant differences between groups for some of the tasks indicates the need for more refined metrics. Although hand-crafted metrics were effective for some simple cases, keeping them explainable limits the complexity they can reach. In effect, not all tasks were able to produce discriminant metrics for all groups of participants, and not all expected behaviors usually observed in handwriting were observed through the proposed handcrafted metrics (for example, micrographia [11] was not observed in any of the tasks for PD participants). Future work should prioritize developing machine learning models that better capture the variations in the most complex patterns observed in handwriting.
Given the multimodal nature of the dataset presented, the integration of handwriting with synchronized speech and eye-tracking data offers a compelling avenue for enhanced assessment performance, which is out of the scope of this study. Future efforts will explore the potential benefits of integrating all available modalities to enhance performance on the task.
Real-world deployment will require portable, user-friendly tools. This work lays the foundation for handwriting analysis as a clinically impactful method for ND characterization, but presents limits that will be addressed in future works.